### v1

# Supplementary Materials

Supplementary Materials
[…]
Appendix A Details on Data Construction
A.1 Details on Data Collection
Table 1: Opioids and their generic naming that used for filtering.
Medication Names Generic Names
Ascomp with Codeine aspirin/butalbital/caffeine/codeine
B & O Supprettes belladonna/opium
Darvon Compound-65 aspirin/caffeine/propoxyphene
Lorcet acetaminophen/hydrocodone
Maxidone acetaminophen/hydrocodone
Fiorinal with Codeine III aspirin/butalbital/caffeine/codeine
Magnacet acetaminophen/oxycodone
Meprozine meperidine/promethazine
Fiorinal with Codeine aspirin/butalbital/caffeine/codeine
Fioricet with Codeine acetaminophen/butalbital/caffeine/codeine
Lorcet Plus acetaminophen/hydrocodone
Percocet 10 / 325 acetaminophen/oxycodone
Primlev acetaminophen/oxycodone
Suboxone buprenorphine/naloxone
Ibudone hydrocodone/ibuprofen
Lorcet 10 / 650 acetaminophen/hydrocodone
Panlor DC acetaminophen/caffeine/dihydrocodeine
Reprexain hydrocodone/ibuprofen
Percocet acetaminophen/oxycodone
Combunox ibuprofen/oxycodone
Hydrocet acetaminophen/hydrocodone
Roxicet acetaminophen/oxycodone
Tylox acetaminophen/oxycodone
Xolox acetaminophen/oxycodone
Vicodin ES acetaminophen/hydrocodone
Hycet acetaminophen/hydrocodone
Talacen acetaminophen/pentazocine
Vicodin HP acetaminophen/hydrocodone
Vicoprofen hydrocodone/ibuprofen
Percocet 7.5 / 325 acetaminophen/oxycodone
Lortab acetaminophen/hydrocodone
Norco acetaminophen/hydrocodone
Vicodin acetaminophen/hydrocodone
Percocet 5 / 325 acetaminophen/oxycodone
Stagesic acetaminophen/hydrocodone
Targiniq ER naloxone/oxycodone
Xodol acetaminophen/hydrocodone
Endocet acetaminophen/oxycodone
Ultracet acetaminophen/tramadol
Panlor SS acetaminophen/caffeine/dihydrocodeine
Zubsolv buprenorphine/naloxone
Xartemis XR acetaminophen/oxycodone
Talwin Nx naloxone/pentazocine
Tylenol with Codeine acetaminophen/codeine
Anexsia acetaminophen/hydrocodone
Darvocet-N 50 acetaminophen/propoxyphene
Liquicet acetaminophen/hydrocodone
Darvocet-N 100 acetaminophen/propoxyphene
Trezix acetaminophen/caffeine/dihydrocodeine
Percodan aspirin/oxycodone
Darvocet A500 acetaminophen/propoxyphene
Percocet 2.5 / 325 acetaminophen/oxycodone
Balacet acetaminophen/propoxyphene
Aceta w/ Codeine acetaminophen/codeine
Zamicet acetaminophen/hydrocodone
Embeda morphine/naltrexone
Bunavail buprenorphine/naloxone
Tylenol with Codeine #3 acetaminophen/codeine
Narvox acetaminophen/oxycodone
Zydone acetaminophen/hydrocodone
Tylenol with Codeine #4 acetaminophen/codeine
Capital w/ Codeine acetaminophen/codeine
Co-Gesic acetaminophen/hydrocodone
Cocet Plus acetaminophen/codeine
Codrix acetaminophen/codeine
Dolacet acetaminophen/hydrocodone
Dolagesic acetaminophen/hydrocodone
Endodan aspirin/oxycodone
Perloxx acetaminophen/oxycodone
Phrenilin with Caffeine and Codeine acetaminophen/butalbital/caffeine/codeine
Roxilox acetaminophen/oxycodone
Synalgos-DC aspirin/caffeine/dihydrocodeine
Theracodophen Low 90 acetaminophen/hydrocodone
Tramapap acetaminophen/tramadol
Trycet acetaminophen/propoxyphene
Verdrocet acetaminophen/hydrocodone
Zolvit acetaminophen/hydrocodone
Astramorph PF morphine
Ionsys fentanyl
Lazanda fentanyl
Levo-Dromoran levorphanol
Numorphan oxymorphone
Onsolis fentanyl
Oxyfast oxycodone
Palladone hydromorphone
Roxanol morphine
Roxanol-T morphine
Roxicodone Intensol oxycodone
Meperitab meperidine
Methadone Diskets methadone
Actiq fentanyl
Fentora fentanyl
Subutex buprenorphine
Demerol meperidine
Dolophine methadone
Roxicodone oxycodone
Duragesic-25 fentanyl
Infumorph morphine
Methadose methadone
Ultram ODT tramadol
Dilaudid hydromorphone
Subsys fentanyl
MSIR morphine
OxyContin oxycodone
Paregoric opium
Duragesic-100 fentanyl
Abstral fentanyl
Oxydose oxycodone
Stadol butorphanol
Duragesic fentanyl
Duragesic-50 fentanyl
Buprenex buprenorphine
Zohydro ER hydrocodone
Duragesic-75 fentanyl
MS Contin morphine
Kadian morphine
Opana oxymorphone
Opana ER oxymorphone
Sublimaze fentanyl
Exalgo hydromorphone
Opium Deodorized opium
Oxaydo oxycodone
Avinza morphine
Nucynta ER tapentadol
Darvon-N propoxyphene
OxyIR oxycodone
Nubain nalbuphine
Dilaudid-HP hydromorphone
Rybix ODT tramadol
Ultram ER tramadol
Butrans buprenorphine
Darvon propoxyphene
Oramorph SR morphine
Nucynta tapentadol
Ultram tramadol
Duramorph morphine
Ryzolt tramadol
Talwin pentazocine
Duragesic-12 fentanyl
Alfenta alfentanil
ConZip tramadol
Hysingla ER hydrocodone
Belbuca buprenorphine
Dazidox oxycodone
DepoDur morphine liposomal
ETH-Oxydose oxycodone
Oxecta oxycodone
Probuphine buprenorphine
RMS morphine
Sufenta sufentanil
Ultiva remifentanil
Continued on next page
Table 2: ICD 9 and ICD 10 diagnosis codes relevant to OUD. Note that, all of these codes defined by weiss2020hospital.
ICD code ICD Description
ICD 9 diagnosis codes
304 Opioid type dependence, unspecified
304.01 Opioid type dependence, continuous
304.02 Opioid type dependence, episodic
304.03 Opioid type dependence, in remission
304.7 Combinations of opioid type drug with any other drug dependence, unspecified
304.71 Combinations of opioid type drug with any other drug dependence, continuous
304.72 Combinations of opioid type drug with any other drug dependence, episodic
304.73 Combinations of opioid type drug with any other drug dependence, in remission
305.5 Opioid abuse, unspecified
305.51 Opioid abuse, continuous
305.52 Opioid abuse, episodic
305.53 Opioid abuse, in remission
965 Poisoning by opium (alkaloids), unspecified
965.01 Poisoning by heroin
965.02 Poisoning by methadone
965.09 Poisoning by other opiates and related narcotics
970.1 Poisoning by opiate antagonists
E850.0 Accidental poisoning by heroin
E850.1 Accidental poisoning by methadone
E850.2 Accidental poisoning by other opiates and related narcotics
E935.0 Heroin causing adverse effects in therapeutic use
E935.1 Methadone causing adverse effects in therapeutic use
E935.2 Other opiates and related narcotics causing adverse effects in therapeutic use
E940.1 Adverse effects of opiate antagonists
ICD 10 diagnosis codes
Opioid abuse/dependence
F11.10 Opioid abuse, uncomplicated
F11.120 Opioid abuse with intoxication, uncomplicated
F11.121 Opioid abuse with intoxication, delirium
F11.122 Opioid abuse with intoxication, with perceptual disturbance
F11.129 Opioid abuse with intoxication, unspecified
F11.14 Opioid abuse with opioid-induced mood disorder
F11.150 Opioid abuse with opioid-induced psychotic disorder, with delusions
F11.151 Opioid abuse with opioid-induced psychotic disorder, with hallucinations
F11.159 Opioid abuse with opioid-induced psychotic disorder, unspecified
F11.181 Opioid abuse with opioid-induced sexual dysfunction
F11.182 Opioid abuse with opioid-induced sleep disorder
F11.188 Opioid abuse with other opioid-induced disorder
F11.19 Opioid abuse with unspecified opioid-induced disorder
F11.20 Opioid dependence, uncomplicated
F11.21 Opioid dependence, in remission
F11.220 Opioid dependence with intoxication, uncomplicated
F11.221 Opioid dependence with intoxication, delirium
F11.222 Opioid dependence with intoxication, with perceptual disturbance
F11.229 Opioid dependence with intoxication, unspecified
F11.23 Opioid dependence with withdrawal
F11.24 Opioid dependence with opioid-induced mood disorder
F11.250 Opioid dependence with opioid-induced psychotic disorder, with delusions
F11.251 Opioid dependence with opioid-induced psychotic disorder, with hallucinations
F11.259 Opioid dependence with opioid-induced psychotic disorder, unspecified
F11.281 Opioid dependence with opioid-induced sexual dysfunction
F11.282 Opioid dependence with opioid-induced sleep disorder
F11.288 Opioid dependence with other opioid-induced disorder
F11.29 Opioid dependence with unspecified opioid-induced disorder
Opioid use
F11.90 Opioid use, unspecified, uncomplicated
F11.920 Opioid use, unspecified with intoxication, uncomplicated
F11.921 Opioid use, unspecified with intoxication delirium
F11.922 Opioid use, unspecified with intoxication, with perceptual disturbance
F11.929 Opioid use, unspecified with intoxication, unspecified
F11.93 Opioid use, unspecified, with withdrawal
F11.94 Opioid use, unspecified, with opioid-induced mood disorder
F11.950 Opioid use, unspecified with opioid-induced psychotic disorder, with delusions
F11.951 Opioid use, unspecified with opioid-induced psychotic disorder, with hallucinations
F11.959 Opioid use, unspecified with opioid-induced psychotic disorder, unspecified
F11.981 Opioid use, unspecified with opioid-induced sexual dysfunction
F11.982 Opioid use, unspecified with opioid-induced sleep disorder
F11.988 Opioid use, unspecified with other opioid-induced disorder
F11.99 Opioid use, unspecified, with unspecified opioid-induced disorder
Poisoning
T40.0X1A Poisoning by opium, accidental (unintentional), initial encounter
T40.0X1D Poisoning by opium, accidental (unintentional), subsequent encounter
T40.0X2A Poisoning by opium, intentional self-harm, initial encounter
T40.0X2D Poisoning by opium, intentional self-harm, subsequent encounter
T40.0X3A Poisoning by opium, assault, initial encounter
T40.0X3D Poisoning by opium, assault, subsequent encounter
T40.0X4A Poisoning by opium, undetermined, initial encounter
T40.0X4D Poisoning by opium, undetermined, subsequent encounter
T40.1X1A Poisoning by heroin, accidental (unintentional), initial encounter
T40.1X1D Poisoning by heroin, accidental (unintentional), subsequent encounter
T40.1X2A Poisoning by heroin, intentional self-harm, initial encounter
T40.1X2D Poisoning by heroin, intentional self-harm, subsequent encounter
T40.1X3A Poisoning by heroin, assault, initial encounter
T40.1X3D Poisoning by heroin, assault, subsequent encounter
T40.1X4A Poisoning by heroin, undetermined, initial encounter
T40.1X4D Poisoning by heroin, undetermined, subsequent encounter
T40.2X1A Poisoning by other opioids, accidental (unintentional), initial encounter
T40.2X1D Poisoning by other opioids, accidental (unintentional), subsequent encounter
T40.2X2A Poisoning by other opioids, intentional self-harm, initial encounter
T40.2X2D Poisoning by other opioids, intentional self-harm, subsequent encounter
T40.2X3A Poisoning by other opioids, assault, initial encounter
T40.2X3D Poisoning by other opioids, assault, subsequent encounter
T40.2X4A Poisoning by other opioids, undetermined, initial encounter
T40.2X4D Poisoning by other opioids, undetermined, subsequent encounter
T40.3X1A Poisoning by methadone, accidental (unintentional), initial encounter
T40.3X1D Poisoning by methadone, accidental (unintentional), subsequent encounter
T40.3X2A Poisoning by methadone, intentional self-harm, initial encounter
T40.3X2D Poisoning by methadone, intentional self-harm, subsequent encounter
T40.3X3A Poisoning by methadone, assault, initial encounter
T40.3X3D Poisoning by methadone, assault, subsequent encounter
T40.3X4A Poisoning by methadone, undetermined, initial encounter
T40.3X4D Poisoning by methadone, undetermined, subsequent encounter
T40.4X1A Poisoning by synthetic narcotics, accidental (unintentional), initial encounter
T40.4X1D Poisoning by synthetic narcotics, accidental (unintentional), subsequent encounter
T40.4X2A Poisoning by other synthetic narcotics, intentional self-harm, initial encounter
T40.4X2D Poisoning by other synthetic narcotics, intentional self-harm, subsequent encounter
T40.4X3A Poisoning by other synthetic narcotics, assault, initial encounter
T40.4X3D Poisoning by other synthetic narcotics, assault, subsequent encounter
T40.4X4A Poisoning by synthetic narcotics, undetermined, initial encounter
T40.4X4D Poisoning by synthetic narcotics, undetermined, subsequent encounter
T40.601A Poisoning by unspecified narcotics, accidental (unintentional), initial encounter
T40.601D Poisoning by unspecified narcotics, accidental (unintentional), subsequent encounter
T40.602A Poisoning by unspecified narcotics, intentional self-harm, initial encounter
T40.602D Poisoning by unspecified narcotics, intentional self-harm, subsequent encounter
T40.603A Poisoning by unspecified narcotics, assault, initial encounter
T40.603D Poisoning by unspecified narcotics, assault, subsequent encounter
T40.604A Poisoning by unspecified narcotics, undetermined, initial encounter
T40.604D Poisoning by unspecified narcotics, undetermined, subsequent encounter
T40.691A Poisoning by other narcotics, accidental (unintentional), initial encounter
T40.691D Poisoning by other narcotics, accidental (unintentional), subsequent encounter
T40.692A Poisoning by other narcotics, intentional self-harm, initial encounter
T40.692D Poisoning by other narcotics, intentional self-harm, subsequent encounter
T40.693A Poisoning by other narcotics, assault, initial encounter
T40.693D Poisoning by other narcotics, assault, subsequent encounter
T40.694A Poisoning by other narcotics, undetermined, initial encounter
T40.694D Poisoning by other narcotics, undetermined, subsequent encounter
Adverse effects
T40.0X5A Adverse effect of opium, initial encounter
T40.0X5D Adverse effect of opium, subsequent encounter
T40.2X5A Adverse effect of other opioids, initial encounter
T40.2X5D Adverse effect of other opioids, subsequent encounter
T40.3X5A Adverse effect of methadone, initial encounter
T40.3X5D Adverse effect of methadone, subsequent encounter
T40.4X5A Adverse effect of synthetic narcotics, initial encounter
T40.4X5D Adverse effect of synthetic narcotic, subsequent encounter
T40.605A Adverse effect of unspecified narcotics, initial encounter
T40.605D Adverse effect of unspecified narcotics, subsequent encounter
T40.695A Adverse effect of other narcotics, initial encounter
T40.695D Adverse effect of other narcotics, subsequent encounter
Long-term use of opiates
Z79.891 Long-term (current) use of opiate analgesic
Continued on next page
A.2 Annotation Guideline
Confirmed aberrant behavior (CAB):
This class refers to behavior that more likely lead to a catastrophic adverse events. It is defined as evidence confirming loss of control of opioid use, specifically aberrant usage of opioid medications, including: 1) Aberrant use of opioids, such as administration/consumption in a way other than described or self-escalating doses. 2) Evidence suggesting or proving that patient has been selling or giving away opioids to others, including family members. 3) Use of opioids for a different indication other than the indication intended by the prescriber. 4) Phrases suggesting current use of illicit or illicitly obtained substances or misuse of legal substances (e.g. alcohol) other than prescription opioid medications.
Suggested aberrant behavior (SAB):
This class refers to behavior implying patient distress related to their opioid treatment. SAB includes three kinds of behavior that suggest potential misuse of opioid. 1) Patient attempt to get extra opioid medicine like requesting for early refill, asking for increasing dosage or reporting missing/stolen opioid medication. 2) Patient emotions toward opioid like request of a certain opioid medication use/change/increase. 3) Physician concerns.
Opioid medication names:
This class refers to the mention or listing of the name(s) of the opioid medication(s) that the patient is currently prescribed or has just been newly prescribed.
Indicated use of the opioid:
This class indicates that patients are using opioid under instructions, such as using opioid for pain, for treatment of opioid use disorder, etc.
Diagnoses of opioid dependence:
It refers to patients have the condition of being dependent on opioids, have chronic opioid use, or is undergoing opioid titration.
Benzodiazepines prescription:
This class refers co-prescribed benzodiazepines (a risk factor for accidental opioid overdose [sun2017association]). In this case, the patient is simply being co-prescribed benzodiazepines (with no noted evidence for abuse).
Medicine changes in an opioid prescription:
This class indicates that the physician makes changes to the patient’s opioid regimen during this current encounter or the patient’s opioid regimen has been changed since the patient’s last encounter with the provider writing the note.
CNS-related:
This is defined as CNS-related terms or terms suggesting altered sensorium, including cognitive impairment, sedation, lightheadedness, intoxication and general term suggesting altered sensorium (e.g. “altered mental status”).
Social Determinants of Health:
This class refers to the factors in the surroundings which impact their well-being. Our dataset captured following attributes:
  • Marital status (single, married …)
  • Cohabitation status (live alone, lives with others …)
  • Educational level (graduate degree, college degree, high-school diploma …)
  • Socioeconomic status (retired, disabled, pension, working …)
  • Homelessness (past, present …)
A.3 Annotation Environment
The annotators coded the EHR with the eHOST annotation tool. Herein,
Appendix B Experimental Setting
B.1 Hyper-parameter Setting
B.2 Language Model Selection